OpenStudy (anonymous):

Solve the compound inequality and answer in interval notation: -6x+5>=23 or 3x+3>=-12

OpenStudy (anonymous):

-6x+5>=23 -6x >= 18 -x >= 3 x <= -3 3x+3>=-12 3x >= -15 x >= -5 so -> -5 <= x <= -3

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

Thanks guys :) That's what I thought but I wasn't sure